# Adaptive Accelerated Composite Minimization

### v1

@@ 3. Accelerated Adaptive Stepsize
This section pertains to the analysis of convergence for update direction $d_{k}$ in accelerated methods. The general form of our accelerated minimization algorithm is provided in Alg. 2, which follows the accelerated proximal gradient descent method with the difference in the choice of stepsize $\lambda$ [29, 32, 12]. Differently from [29, 8], the stepsize is not fixed, and in our numerical experience, it does not strictly decrease in many iterations, which partially explains the acceleration of the algorithm up to the optimal theoretical bound $\mathcal{O}(k^{-2})$.
@@
$\begin{cases}\beta_{k}=\dfrac{1+\sqrt{1+4\beta_{k-1}^{2}}}{2},\,\,\gamma_{k}=\dfrac{1-\beta_{k}}{\beta_{k+1}},\,\,x_{0}=x_{1},\,\,\beta_{1}=0,\\ \lambda_{k}=\arg\max\limits_{\lambda\in\mathbb{R}}\left\{\lambda\,|\,\,\phi(2\lambda)\leq\phi(\lambda)-\lambda\langle G^{f}_{\lambda h}(x),\nabla f(x)\rangle+\dfrac{\lambda}{2}\|G^{f}_{\lambda h}(x))\|^{2},\,\lambda\leq\lambda_{k-1}\right\},\\ d_{k}=-\gamma_{k}\left(\dfrac{\lambda_{k-1}}{\lambda_{k}}d_{k-1}-(1+\dfrac{1}{\gamma_{k}})G^{f}_{\lambda_{k}h}(x_{k})+\dfrac{\lambda_{k-1}}{\lambda_{k}}\nabla G^{f}_{\lambda_{k-1}h}(x_{k-1})\right).\end{cases}$ (Alg. 2)
ensures the uniform stepsize lower bound $\lambda_{k}\geq{1}/{2L}$ and the sub-optimality bound $F(x_{k})-F(x^{*})\leq\dfrac{D^{\prime}}{k^{2}}$ where $L$ is the smoothness constant of $f$, and $D^{\prime}$ is a constant that only depends on the initial condition $x_{0}$, the optimal solution $x^{*}$, and the constant $L$.
@@
The proof of a first proposition is the same as in Theorem 2.3. Specifically, according to stepsize rule in Alg. 2, we should first check whether the previous $\lambda_{k-1}$ satisfies the linesearch condition. Since the inequality (14) holds for $\lambda={1}/{2L}$ (according to the linesearch condition in Alg. 2), we have the guarantee that $\lambda_{k}\geq{1}/{2L}$. Regarding the convergence analysis bound, in the first step, we invoke the increment inequality of (iii) in Lemma 2.2 for two cases of $z=y_{k}$ and $z=x^{*}$, which arrives at
@@
$\|\beta_{k}(y_{k+1}-x_{k})\|^{2}+2\beta_{k}\langle y_{k+1}-x_{k},\beta_{k}x_{k}-(\beta_{k}-1)y_{k}-x^{*}\rangle=$
@@
$\lambda_{T}\beta_{T}^{2}\delta_{T+1}-\lambda_{0}\beta_{0}^{2}\delta_{1}\leq\dfrac{1}{2}\|u_{1}\|^{2}-\dfrac{1}{2}\|u_{T+1}\|^{2}\leq\dfrac{1}{2}||u_{1}||^{2}.$
@@
$F(y_{T+1})-F(x^{*})=\delta_{T+1}\leq\dfrac{4L\|u_{1}\|^{2}}{CT^{2}}\leq\dfrac{D^{\prime}}{T^{2}}.$
@@
Let the function $F$ in (6) be a smooth (i.e., $h(x)=0$). Then, the stepsize rule in Alg. 2 reduces to (7) while ensuring Theorem 3.1 results of the uniform stepsize lower bound $\lambda_{k}\geq{1}/{2L}$ and the sub-optimality bound $F(x_{k})-F(x^{*})\leq\dfrac{D^{\prime}}{k^{2}}$.
@@ 4. Numerical Results
Let us first start the numerical section with a practical discussion concerning the implementation of the approximate solution in Remark 2.5 for the stepsize rule in Alg. 1. The backtracking in Remark 2.5 uses an initial constant $\lambda_{0}$ to approximate the proposed stepsize rule in Alg. 1. On the one hand, if the initial $\lambda_{0}$ is too large, it takes time to compute the desired stepsize satisfying linesearch condition. On the other hand, a too small $\lambda_{0}$ deteriorates the convergence speed [34]. To address this issue, we approximate the function $\phi$ in (2) via a quadratic function, denoted by $\widetilde{\phi}$, and suggest the desired $\lambda_{0}$ as its optimizer; see Figure 2 for a visual representation of this approximation (cf. Figure 1b).
@@
We illustrate the performance of our proposed stepsize rules Alg. 1 and Alg. 2 in three cases of (1) smooth minimization including (1-i) logistic regression, (1-ii) quadratic programming, (1-iii) log-sum-exp, and (1-iv) max-cut as a representative case of approximate semidefinite programming, (2) composite minimization including (2-i) $\ell_{1}$-Regularized least square, (2-ii) $\ell_{1}$-Constrained least square, (2-iii) $\ell_{1}$-Regularized logistic regression, and (3) non-convex minimization where we consider a cubic objective function. To evaluate our performance, we compare our proposed algorithms with the following methods from the literature: For the class of (1) smooth optimization, we consider the gradient descent (GD) [34], Nesterov’s accelerated gradient descent (NAGD) [29], the recent adaptive gradient descent (AdGD) and its heuristic accelerated version (AdGD-accel) [24]. Respectively, for the class of (2) composite minimization, we consider the nonsmooth counterparts including the proximal gradient descent (ISTA) and its accelerated version (FISTA) [8], the adaptive golden ratio algorithm (aGRAAL) [23] and adaptive primal-dual algorithm (APDA) [39]. For (3) non-convex minimization, we use the existing algorithms from the first class of (1) smooth minimization.
@@ 4.1. Smooth minimization
(1-i) Logistic regression
Consider the logistic regression problem with quadratic regularization
@@
where $N$ is a number of data, $A=[a_{1}|a_{2}|\cdots|a_{n}]\in\mathbb{R}^{d\times N}$ and $\{b_{i}\}_{i=1}^{N}\in\mathbb{R}$ are the features of data, and $\gamma$ is a regularization parameter proportional to ${1}/{N}$. In our experiments, The test data is generated as follows: the matrix $A\in\mathbb{R}^{N\times N}$ is generated randomly using the standard Gaussian distribution $\mathcal{N}(0,1)$, where we generate $A$ with $50\%$ correlated columns as $A(:,j+1)=0.5A(:,j)+\textrm{randn}(:)$. The observed measurement vector $b$ is generated as $b:=Ax^{\natural}+\mathcal{N}(0,0.05)$, where $x^{\natural}$ is generated randomly using $\mathcal{N}(0,1)$, $d=N=200$, and we estimate the Lipschitz constant of the gradient by the closed-form expression $L=\|A\|^{2}/N+\gamma$ used in GD and NAGD as a stepsize. Figure 3a reports the results where all algorithms are initialized at the same point chosen randomly.
(1-ii) Quadratic programming
Consider the quadratic objective function
@@
(1-iii) Log-Sum-Exp
Consider the log-sum-exp objective function with regularization
@@
where $A=[a_{1}|a_{2}|\cdots|a_{n}]\in\mathbb{R}^{d\times N}$ and $\{b_{i}\}_{i=1}^{N}\in\mathbb{R}$ are the problem data. The log-sum-exp function is viewed as a smooth approximation of $\max_{x\in\mathbb{R}^{d}}\{a_{1}^{\top}x-b_{1},a_{2}^{\top}x-b_{2},\cdots,a_{N}^{\top}x-b_{N}\}$. This function can sometimes be ill-conditioned, making the minimization task particularly difficult for the first-order methods. The function is smooth with the smoothness constant $L=\sigma_{max}(A^{\top})$ where $\sigma_{max}(A^{\top})$ is a maximum singular value of $A^{\top}$. In the experiment, we set $N=d=200$, and $A$ and $\{b_{i}\}_{i=1}^{N}$ are generated using a similar approach as in the logistic regression part. The results of minimizing the log-sum-exp function are depicted in Figure 3c where the proposed adaptive accelerated method demonstrates superior efficacy compared to other algorithms.
@@
(1-vi) Approximate semidefinite programming: the case of max-cut
Semidefinite programs are ubiquitous in control problems [11], graph theory [38], network and communication [9]. Most of the problems in this class essentially deal with minimizing the maximum eigenvalue of a matrix, which is a nonsmooth function of its variable. A popular example falling into this category is the max-cut problem, whose $\varepsilon$-approximation with the regularizer $\mathcal{R}(x)$ reads as
@@
where $C$ is a constant matrix, $\eta$ is a regularization coefficient, and $f_{\varepsilon}(X)$ is the smooth convex approximation of $\lambda_{\textbf{max}}(X)$. As shown in [30], the gradient of $f_{\varepsilon}(X)$ can be computed as
@@
where $q_{i}$ is the $i^{\text{th}}$ column of the unitary matrix $Q$ in the eigen-decomposition $Q\Sigma Q^{\top}$ of $X$ with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{n}$. In addition, $f_{\varepsilon}(X)$ is smooth with the smoothness parameter $1/\varepsilon$. Figure 4 shows the simulation results of solving (28) for different regularization coefficient $\eta$ where the approximation level is $\varepsilon=10^{-5}$. The matrix $C$ is also generated using the Wishart distribution with $C={G^{\top}G}/{\|G\|_{2}^{2}}$, where $G$ is a standard Gaussian matrix [17], $n=100$, and $\mathcal{R}(y)=\|y\|^{2}$.
@@ 4.2. Composite minimization
For the case of composite minimization, we consider the following three nonsmooth functions:
(2-i) $\ell_{1}$-Regularized least square
$\min\limits_{x\in\mathbb{R}^{d}}\underbrace{\|Ax-b\|_{2}^{2}}_{f(x)}+\underbrace{\gamma\|x\|_{1}}_{h(x)}$ (29)
(2-ii) $\ell_{1}$-Constrained least square
$\min\limits_{x\in\mathbb{R}^{d}}\{\|Ax-b\|_{2}^{2}\,:\|x\|_{1}\leq 1\}=\min\limits_{x\in\mathbb{R}^{d}}\underbrace{\|Ax-b\|_{2}^{2}}_{f(x)}+\underbrace{\delta_{B_{1}[0,1]}}_{h(x)}$ (30)
(2-iii) $\ell_{1}$-Regularized logistic regression
$\min\limits_{x\in\mathbb{R}^{d}}\underbrace{\dfrac{1}{N}\sum_{i=1}^{N}\log\left(1+\exp(-b_{i}a_{i}^{\top}x)\right)}_{f(x)}+\underbrace{\gamma\|x\|_{1}}_{h(x)}$ (31)
The functions $f(x)$ and $h(x)$ represent the smooth and prox-friendly parts, respectively. In the previous subsection, we explain the smooth parts and their smooth constant. The prox-friendly term $h(x)$ has a closed form solution provided in Table 2 where $\delta$, $B_{1}[0,1]$, $[\cdot]_{+}$, and $\odot$ are indicator function, unit ball with 1-norm, positive part selector, and elementwise product, respectively [6].
@@
[…]
@@
4.3. Non-convex minimization
We also test our algorithm in non-convex minimization. We consider the case of cubic optimization that is useful in high-order methods [33] and is defined as
$\min_{x\in\mathbb{R}^{d}}\dfrac{1}{2}x^{\top}Hx+g^{\top}x+\dfrac{M}{6}\|x\|^{3}.$ (32)
The matrcies $H\in\mathbb{R}^{d\times d}$, $g\in\mathbb{R}^{d}$, and $M>0$ are the problem data. Note that in higher-order methods, the constant $M$ has an impact on the convergence rate [31]. Due to the cubic term, the function (32) is neither convex nor smooth. The parameters $g$ and $H$ are the gradient and the Hessian of the logistic loss (25), respectively, and are computed using the same data in part (1-i) logistic regression. We provide the simulations for different $M$ with $d=N=200$, and the results are reported in Figure 6.

### v2

@@ 3. Accelerated Adaptive Stepsize
This section pertains to the analysis of convergence for update direction $d_{k}$ in accelerated methods. The general form of our accelerated minimization algorithm is provided in Alg. 2, which follows the accelerated proximal gradient descent method with the difference in the choice of stepsize $\lambda$ [Nesterov1983, Nesterov2018, bubeck2015convex]. Differently from [Nesterov1983, Beck2009], the stepsize is not fixed, and in our numerical experience, it does not strictly decrease in many iterations, which partially explains the acceleration of the algorithm up to the optimal theoretical bound $\mathcal{O}(k^{-2})$.
@@
$\begin{cases}\beta_{k}=\dfrac{1+\sqrt{1+4\beta_{k-1}^{2}}}{2},\,\,\gamma_{k}=\dfrac{1-\beta_{k}}{\beta_{k+1}},\,\,x_{0}=x_{1},\,\,\beta_{1}=0,\\ \lambda_{k}=\arg\max\limits_{\lambda\in\mathbb{R}}\left\{\lambda\,|\,\,\phi(2\lambda)\leq\phi(\lambda)-\lambda\langle G^{f}_{\lambda h}(x),\nabla f(x)\rangle+\dfrac{\lambda}{2}\|G^{f}_{\lambda h}(x))\|^{2},\,\lambda\leq\lambda_{k-1}\right\},\\ d_{k}=-\gamma_{k}\left(\dfrac{\lambda_{k-1}}{\lambda_{k}}d_{k-1}-(1+\dfrac{1}{\gamma_{k}})G^{f}_{\lambda_{k}h}(x_{k})+\dfrac{\lambda_{k-1}}{\lambda_{k}}G^{f}_{\lambda_{k-1}h}(x_{k-1})\right).\end{cases}$ (Alg. 2)
ensures the uniform stepsize lower bound $\lambda_{k}\geq{1}/{3L}$ and the sub-optimality bound $F(x_{k})-F(x^{*})\leq\dfrac{D^{\prime}}{k^{2}}$ where $L$ is the smoothness constant of $f$, and $D^{\prime}$ is a constant that only depends on the initial condition $x_{0}$, the optimal solution $x^{*}$, and the constant $L$.
@@
The proof of a first proposition is the same as in Theorem 2.3. Specifically, according to stepsize rule in Alg. 2, we should first check whether the previous $\lambda_{k-1}$ satisfies the linesearch condition. Since the inequality (14) holds for $\lambda={1}/{3L}$ (according to the linesearch condition in Alg. 2), we have the guarantee that $\lambda_{k}\geq{1}/{3L}$. Regarding the convergence analysis bound, in the first step, we invoke the increment inequality of (iii) in Lemma 2.2 for two cases of $z=y_{k}$ and $z=x^{*}$, which arrives at
@@
[…]
@@
[…]
@@
$F(y_{T+1})-F(x^{*})=\delta_{T+1}\leq\dfrac{6L\|u_{1}\|^{2}}{CT^{2}}\leq\dfrac{D^{\prime}}{T^{2}}.$
@@
Let the function $F$ in (6) be a smooth (i.e., $h(x)=0$). Then, the stepsize rule in Alg. 2 reduces to (7) while ensuring Theorem 3.1 results of the uniform stepsize lower bound $\lambda_{k}\geq{1}/{3L}$ and the sub-optimality bound $F(x_{k})-F(x^{*})\leq\dfrac{D^{\prime}}{k^{2}}$.
@@ 4. Numerical Results
We illustrate the performance of our proposed stepsize rules Alg. 1 and Alg. 2 in three major problem classes widely used in operations research, machine learning, and systems and control:
  1. (1)
    Smooth functions, including the specific objective functions
    1. (i)
      logistic regression, motivated by classification tasks in machine learning problems [jiang2020online],
    2. (ii)
      quadratic programming, motivated by portfolio optimization and resource allocation problems [huang2025restarted, mittal2020robust, mehrotra1990algorithm]),
    3. (iii)
      log-sum-exp, motivated by discrete-choice modeling and entropy-regularized transport, e.g., multinomial logit model and softmax demand models [feng65technical, qu2025sinkhorn, liu2020assortment], and
    4. (iv)
      max-cut, a benchmark for combinatorial optimization and approximate semidefinite programming [boyd1997semidefinite, nesterov2007smoothing, han2025low].
  2. (2)
    Composite minimization, including the specific objective functions
    1. (i)
      $\ell_{1}$-regularized least squares, commonly used in high-dimensional statistics for support recovery and in signal processing for sparse signal reconstruction [neykov2016l1, selesnick2017sparse]),
    2. (ii)
      $\ell_{1}$-constrained least squares, in sparse signal recovery with explicit coefficient bounds [gaines2018algorithms]), and (2-iii) $\ell_{1}$-regularized logistic regression (sparse classification and feature selection, [bertsimas2021sparse]).
  3. (3)
    Non-convex minimization, where we consider a cubic objective function relevant to cubic-regularization in nonconvex optimization and trust-region subproblems [Nesterov2006, chen2022accelerating].
Before presenting the numerical results, let us first start with a practical discussion concerning the implementation of the approximate solution in Remark 2.5 for the stepsize rule in Alg. 1. The backtracking in Remark 2.5 uses an initial constant $\lambda_{0}$ to approximate the proposed stepsize rule in Alg. 1. On the one hand, if the initial $\lambda_{0}$ is too large, it takes time to compute the desired stepsize satisfying linesearch condition. On the other hand, a too small $\lambda_{0}$ deteriorates the convergence speed [Nocedal2006]. To address this issue, we approximate the function $\phi$ in (2) via a quadratic function, denoted by $\widetilde{\phi}$, and suggest the desired $\lambda_{0}$ as its optimizer; see Figure 2 for a visual representation of this approximation (cf. Figure 1b).
@@
To evaluate our performance, we compare our proposed algorithms with the following methods from the literature: For the class of (1) smooth optimization, we consider the gradient descent (GD) [Nocedal2006], Nesterov’s accelerated gradient descent (NAGD) [Nesterov1983], the recent adaptive gradient descent (AdGD) and its heuristic accelerated version (AdGD-accel) [Malitsky2020]. Respectively, for the class of (2) composite minimization, we consider the nonsmooth counterparts including the proximal gradient descent (ISTA) and its accelerated version (FISTA) [Beck2009], the adaptive golden ratio algorithm (aGRAAL) [malitsky2020golden] and adaptive primal-dual algorithm (APDA) [vladarean2021first]. For (3) non-convex minimization, we use the existing algorithms from the first class of (1) smooth minimization.
@@ 4.1. Smooth minimization
(i) Logistic regression:
Motivating with the classification problems [jiang2020online], we consider the logistic regression with quadratic regularization as follows:
@@
where $N$ is a number of data, $A=[a_{1}|a_{2}|\cdots|a_{n}]\in\mathbb{R}^{d\times N}$ and $\{b_{i}\}_{i=1}^{N}\in\mathbb{R}$ are the features of data, and $\gamma$ is a regularization parameter proportional to 1/N{1}/{N}. In our experiments, The test data is generated as follows: the matrix $A\in\mathbb{R}^{N\times N}$ is generated randomly using the standard Gaussian distribution $\mathcal{N}(0,1)$, where we generate $A$ with $50\%$ correlated columns as $A(:,j+1)=0.5A(:,j)+\textrm{randn}(:)$. The observed measurement vector $b$ is generated as $b:=Ax^{\natural}+\mathcal{N}(0,0.05)$, where $x^{\natural}$ is generated randomly using $\mathcal{N}(0,1)$, $d=N=200$, and we estimate the Lipschitz constant of the gradient by the closed-form expression $L=\|A\|^{2}/N+\gamma$ used in GD and NAGD as a stepsize. Figure 3a reports the results where all algorithms are initialized at the same point chosen randomly.
(ii) Quadratic programming:
Following recent works on portfolio and robust optimization [huang2025restarted, mittal2020robust], we examine the quadratic objective function
@@
(iii) Log-Sum-Exp:
In the context of discrete-choice modeling and entropy-regularized transport [feng65technical, qu2025sinkhorn, liu2020assortment], we consider the log-sum-exp objective function with regularization
@@
where $A=[a_{1}|a_{2}|\cdots|a_{n}]\in\mathbb{R}^{d\times N}$ and $\{b_{i}\}_{i=1}^{N}\in\mathbb{R}$ are the problem data. The log-sum-exp function is viewed as a smooth approximation of $\max_{x\in\mathbb{R}^{d}}\{a_{1}^{\top}x-b_{1},a_{2}^{\top}x-b_{2},\cdots,a_{N}^{\top}x-b_{N}\}$. This function can sometimes be ill-conditioned, making the minimization task particularly difficult for the first-order methods. The function is smooth with the smoothness constant $L=\sigma_{max}(A^{\top})$ where $\sigma_{max}(A^{\top})$ is a maximum singular value of $A^{\top}$. In the experiment, we set $N=d=200$, and $A$ and $\{b_{i}\}_{i=1}^{N}$ are generated using a similar approach as in the logistic regression part. The results of minimizing the log-sum-exp function are depicted in Figure 3c where the proposed accelerated method demonstrates superior efficacy compared to other algorithms.
Refer to captionRefer to caption
@@
(vi) Approximate semidefinite programming:
Semidefinite programs are ubiquitous in combinatorial optimization and control [boyd1997semidefinite, El-Ghaoui, han2025low], graph theory [tang2024solving], network and communication [biswas2006semidefinite]. Most of the problems in this class essentially deal with minimizing the maximum eigenvalue of a matrix, which is a nonsmooth function of its variable. A popular example falling into this category is the max-cut problem, whose $\varepsilon$-approximation with the regularizer $\mathcal{R}(x)$ reads as
@@
where $C$ is a constant matrix, $\eta$ is a regularization coefficient, and $f_{\varepsilon}(X)$ is the smooth convex approximation of $\lambda_{\textbf{max}}(X)$. As shown in [nesterov2007smoothing], the gradient of $f_{\varepsilon}(X)$ can be computed as
@@
where $q_{i}$ is the $i^{\text{th}}$ column of the unitary matrix $Q$ in the eigen-decomposition $Q\Sigma Q^{\top}$ of $X$ with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{n}$. In addition, $f_{\varepsilon}(X)$ is smooth with the smoothness parameter $1/\varepsilon$. Figure 4 shows the simulation results of solving (28) for different regularization coefficient $\eta$ where the approximation level is $\varepsilon=10^{-5}$. The matrix $C$ is also generated using the Wishart distribution with $C={G^{\top}G}/{\|G\|_{2}^{2}}$, where $G$ is a standard Gaussian matrix [helmberg2000spectral], $n=100$, and $\mathcal{R}(y)=\|y\|^{2}$.
@@ 4.2. Composite minimization
Motivated by sparse recovery and reconstruction problems [neykov2016l1, selesnick2017sparse, gaines2018algorithms], as well as sparse classification and feature selection [bertsimas2021sparse], we consider the following non-smooth composite minimization formulations:
(i) $\ell_{1}$-Regularized least square [neykov2016l1, selesnick2017sparse]:
$\min\limits_{x\in\mathbb{R}^{d}}\underbrace{\|Ax-b\|_{2}^{2}}_{f(x)}+\underbrace{\gamma\|x\|_{1}}_{h(x)}.$ (29)
(ii) $\ell_{1}$-Constrained least square [gaines2018algorithms]:
$\min\limits_{x\in\mathbb{R}^{d}}\{\|Ax-b\|_{2}^{2}\,:\|x\|_{1}\leq 1\}=\min\limits_{x\in\mathbb{R}^{d}}\underbrace{\|Ax-b\|_{2}^{2}}_{f(x)}+\underbrace{\delta_{B_{1}[0,1]}}_{h(x)}.$ (30)
(iii) $\ell_{1}$-Regularized logistic regression [bertsimas2021sparse]:
$\min\limits_{x\in\mathbb{R}^{d}}\underbrace{\dfrac{1}{N}\sum_{i=1}^{N}\log\left(1+\exp(-b_{i}a_{i}^{\top}x)\right)}_{f(x)}+\underbrace{\gamma\|x\|_{1}}_{h(x)}.$ (31)
The functions $f(x)$ and $h(x)$ represent the smooth and prox-friendly parts, respectively. In the previous subsection, we explain the smooth parts and their smooth constant. The prox-friendly term $h(x)$ has a closed form solution provided in Table 2 where $\delta$, $B_{1}[0,1]$, $\cdot$$_{+}$, and $\odot$ are indicator function, unit ball with 1-norm, positive part selector, and elementwise product, respectively [Beck2017].
@@
In our simulations, we set $N=d=200$, $\gamma$ proportional to 1/N{1}/{N}, $A=5\cdot\texttt{rand(n,n)}$, and $b=\texttt{rand(n,1)}$, where rand($\cdot$) is uniformly distributed random numbers in the interval $(0,1)$. The results are provided in Figure 5. We wish to note that for the objective functions (29) and (30), the proposed accelerated stepsize Alg. 2 has a competitive result and similar behavior with the best performance FISTA [Beck2009] while in the case of (31) outperforms all the other algorithms notably.
@@
4.3. Non-convex minimization:
We also test our algorithm in non-convex minimization. We consider the case of cubic optimization that is useful in high-order methods and trust-region subproblems [Nesterov2006, chen2022accelerating] which is defined as
$\min_{x\in\mathbb{R}^{d}}\dfrac{1}{2}x^{\top}Hx+g^{\top}x+\dfrac{M}{6}\|x\|^{3},$ (32)
where the matrix $H\in\mathbb{R}^{d\times d}$, vector $g\in\mathbb{R}^{d}$, and scalar $M>0$ are the problem data. Note that in higher-order methods, the constant $M$ has an impact on the convergence rate [Nesterov2008]. Due to the cubic term, the function (32) is neither convex nor smooth. The parameters $g$ and $H$ are the gradient and the Hessian of the logistic loss (25), respectively, and are computed using the same data as in Subsection 4.1 ((i) logistic regression). We provide the simulations for different $M$ with $d=N=200$, and the results are reported in Figure 6.